# Structured Initialization for Vision Transformers

### v1

@@ 1 Introduction
To bridge this performance gap, several strategies have been proposed. These include self-supervised pretraining on large-scale datasets [8, 26], advanced data augmentation techniques [36, 6], and hybrid architectures that incorporate convolutional layers into Vision Transformers [32, 16, 35, 15, 7]. More recently, Zhang et al. [37] explored the use of pretrained weights to initialize ViTs. Building on this idea, subsequent works have shown that carefully designed network initialization strategies can enhance ViT performance on small-scale datasets without modifying the model architecture. In particular, Trockman and Kolter [28] introduced mimetic initialization that replicates the weight distribution of pretrained ViTs. Similarly, Xu et al. [33] proposed directly sampling weights from large pretrained models.
@@
To overcome the limitations of existing approaches, we propose a novel “structured initialization” strategy for ViTs that imposes convolutional structures on attention maps without requiring any pretrained models. Our approach is grounded in our theoretical insights, aligned with the findings of recent work [3], that randomly initialized depthwise convolution filters can match the performance of their trained counterparts in models such as ConvMixer [27] and ResNet [11]. Inspired by this, we develop an initialization strategy for ViTs based on random impulse convolution kernels, which impart locality and structure directly into the attention maps, as shown in Fig. 1. This structured initialization yields inductive biases characteristic of CNNs, enabling ViTs to generalize effectively on small-scale datasets, while preserving their adaptability for large-scale applications. Unlike prior methods that modify ViT architectures by integrating convolutional components, our method preserves the original transformer structure, making it broadly applicable across a variety of ViT variants.
@@ 2 Related Work
Many efforts have aimed to incorporate a convolutional inductive bias into ViTs through architectural modifications. [7] proposed to combine convolution and self-attention by mixing the convolutional self-attention layers. [21, 15]introduced hybrid models wherein the output of each layer is a summation of convolution and self-attention. [32] explored using convolution for token projections within self-attention, while [35] showed promising results by inserting a depthwise convolution before the self-attention map. [9] introduced gated positional self-attention to imply a soft convolution inductive bias. Although these techniques have been proven effective, they aim to introduce the inductive bias of convolution through architectural choices. Our approach, on the other hand, stands out by not requiring any modifications to the architecture, retaining the generalizability to be seamlessly applied to different settings.
@@
The exploration of applying inductive bias through initialization within a transformer is limited to date. [37] posited that the benefit of pretrained models in ViTs can be interpreted as a more effective strategy for initialization. [28, 29] recently investigated the empirical distributions of self-attention weights, learned from large-scale datasets, and proposed a mimetic initialization strategy. While this approach lies between structured and generative initialization, it relies on the pretraining results of large models. [23, 24, 33, 14] directly sampled weights from pretrained large-scale models as initialization for smaller models. While effective, the sampled weights must follow the distribution of these pretrained weights. A key difference in our approach is that our method does not require offline knowledge of pretrained models. Instead, our initialized structure is derived from a theoretical analysis of convolution layers.
@@
Since their introduction [30, 8], the relationship between transformers and CNNs has been a topic of immense interest. [1] studied the structural similarities between attention and convolution, bridging them into a unified framework. Building on this, [5] demonstrated that self-attention layers can express any convolutional layers through a careful theoretical construction. While these studies highlighted the functional equivalence between self-attention in ViTs and convolutional spatial mixing in CNNs, they did not delve into how the inductive bias of ViTs could be adapted through this theoretical connection. In contrast, our work offers a theoretical insight: a random convolutional impulse filter can be effectively approximated by softmax self-attention.
@@ 3 Why Random Impulse Filters Work?
It is well established that both ConvMixer and ViTs use alternating blocks of spatial and channel mixing, where ViTs replace the spatial convolutions in ConvNets with attention mechanisms. A fundamental difference, except for the receptive field, lies in their parameterization: spatial convolutions with hundreds of channels typically use distinct kernels for each channel, whereas self-attention relies on a shared mechanism with only a limited number ( ${\sim}\,10$) of attention heads. By introducing a key observation that input embeddings are often rank-deficient, we demonstrate that as long as the spatial kernels or attention patterns sufficiently span the kernel space, the uniqueness or repetition of individual kernels becomes less critical. This insight, supported by a recent empirical observation [3], offers a new perspective on the relationship between ViTs and CNNs, motivating the use of impulse-based structures to embed convolutional inductive biases into the attention map initialization.
In recent work [3], Cazenavette et al. demonstrated a remarkable performance of randomly initialized convolution filters in ConvMixer and ResNet when solely learning the channel mixing parameters. However, they failed to offer any insights into the underlying reasons. In this section, we provide a theoretical analysis of how solely learning channel mixing can be sufficient for achieving reasonably good performance. Our theoretical findings are significant as they establish a conceptual link between the architecture of ConvMixer and the initialization of ViT, offering a deeper understanding of desired properties for spatial mixing matrices. Without losing generality, we have omitted activations (e.g., GeLU, ReLU, etc.), bias, batch normalization, and skip connections in our equations for clarity.
@@
Let us define the patch embeddings or intermediate layer outputs in ConvMixer as $\mathbf{X}\,{=}\,[\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{D}]$, where $D$ is the number of channels and $N$ is the number of pixels in the vectorized patch $\mathbf{x}\,{\in}\,\mathbb{R}^{N}$. An interesting observation is the rank (stable rank, defined as $\sum\sigma^{2}\,{/}\,\sigma_{max}^{2}$) of $\mathbf{X}$ is consistently much smaller than the minimum dimension $\mbox{min}(N,D)$ of $\mathbf{X}$, indicating a significant amount of redundancy in patch embeddings or intermediate layer outputs in deep networks. This rank deficiency is common in various deep neural networks [10], especially in ViTs [20, 18, 12].
@@
(M–k Spanned Set): Let $\mathcal{V}\,{=}\,\{\mathbf{v}_{1},\mathbf{v}_{2},\ldots,\mathbf{v}_{N}\}\subset\mathbb{R}^{d}$ be a finite set of vectors. We say that $\mathcal{V}$ is $M$$k$ spanned (on $W$) if there exists a partition of $\mathcal{V}$ into at least $k$ non-overlapping subsets:
@@
such that there exist a $M$–dimensional subspace $W\subset\mathbb{R}^{d}$ in the span of each subset $\mathcal{V}_{i}$
@@
A ConvMixer block $\mathbf{T}$ consists of a spatial mixing layer $\mathbf{T}_{S}(\>\cdot\>;\mathbf{H})$ with convolution filters $\mathbf{H}$ and a channel mixing layer $\mathbf{T}_{C}$. $\mathbf{T}^{\prime}$ is another ConvMixer block composed of $\mathbf{T}^{\prime}_{S}(\>\cdot\>;\mathbf{H}^{\prime})$ and $\mathbf{T}^{\prime}_{C}$. Let $k$ be the rank of input $\mathbf{X}$. If $\mathbf{H}$ is $M$$k$ spanned on $W$, then for any set of filters $\mathbf{H}^{\prime}\subset W$ and any $\mathbf{T}^{\prime}_{C}$, there always exists a $\mathbf{T}_{C}$ such that $\mathbf{T}(\mathbf{X})\,{=}\,\mathbf{T}^{\prime}(\mathbf{X})$.
For simplicity, we include the full proof in the appendix. Note that since $H$ are convolution matrices, their span lies in $\mathbb{R}^{f{\times}f}$ instead of $\mathbb{R}^{N{\times}N}$, where $f$ is the kernel size. In practice, the fact that $D\,{\geq}\,kf^{2}$ indicates that randomly initialized spatial convolution kernels are $f^{2}$$k$ spanned and satisfy Proposition 1, as illustrated in Fig. 2. Consequently, any trained results of $\mathbf{T}^{\prime}_{C}$ and $\mathbf{T}^{\prime}_{S}$ can be achieved by solely training the $\mathbf{T}_{C}$, while keeping a fixed spatial mixing layer $\mathbf{T}_{S}$. Hence, the following corollaries can be obtained, and Corollary 1 explains the phenomenon found in [3], mentioned at the beginning of the section. Corollary 2 inspires our proposed structure initialization for ViTs. The related experimental evidence of these corollaries and our proposition is given in the appendix.
@@
[…]
@@ 4.1 Expected Initialized Attention Map Structure
ConvMixer and ViT share most of the components in their architectures. The gap in their performance on small-scale datasets stems from their architectural choices regarding the spatial mixing matrix. Although depthwise convolution (ConvMixer) and multi-head self-attention (ViT) may appear distinct at first glance, their underlying goal remains the same: to identify spatial patterns indicated by the spatial mixing matrix. As defined in Sec. 3, similar to the spatial mixing step in ConvMixer defined in Eq. 1, the spatial mixing step of multi-head attention can be expressed as
@@
where $d$ represents the feature dimension in each head, typically set to $D{/}h$, with $h$ being the number of heads, and the matrices $\mathbf{M}_{i}$ for multi-head self-attention can be expressed as follows:
@@
It is worth noting that in Eq. 1, the spatial matrices $\mathbf{H}$ are in convolutional structure, resulting in a span of $\mathbb{R}^{f{\times}f}$ instead of $\mathbb{R}^{N{\times}N}$, despite each $\mathbf{H}_{i}$ having a size of $N\,{\times}\,N$. This structural constraint ensures that CNNs focus on local features but struggle to capture long-range dependencies. In contrast, the span of spatial matrices $\mathbf{M}$ in Sec. 4.1 is $\mathbb{R}^{N{\times}N}$, allowing for greater learning capacity without these limitations. However, a randomly initialized $\mathbf{Q}$ and $\mathbf{K}$ contain no structural information, resulting in random matrices as depicted in the bottom left of Fig. 1.
Leveraging this insight, we propose to initialize the attention map for each head in ViT to a convolutional structure as denoted in the bottom right of Fig. 1. Our initialization strategy preserves both the advantage of locality and the capacity to learn long-range information. For clarity and brevity, the following discussions will focus only on one head of multi-head self-attention. Therefore, from Sec. 4.1 and Eq. 1, our structured initialization strategy can be represented as
[…]
@@
Any random convolution filters that satisfy proposition 1 could be a choice of initialization of attention maps to introduce inductive bias. However, random convolution filters $\mathbf{H}$ usually contain both positive and negative values, while the output of the softmax function is always positive, making Eq. 8 unreachable. One straightforward option is to use random positive convolution filters with a normalized sum of one, following the property of softmax. However, this approach often proves inefficient as the patterns are too complicated for a softmax function to handle with $\mathbf{Q}\mathbf{K}$ being of low rank. In [25], the authors found that the softmax attention map serves as a feature selection function, which typically tends to select a single related feature. In convolution filters, this selection can be parameterized as impulse filters. According to Corollary 2, random impulse filters are also $f^{2}$$k$ spanned. In conclusion, when initializing a softmax attention map, the most straightforward and suitable choice is random impulse convolution filters.
@@
With the use of impulse filters and the pseudo input, Eq. 8 becomes
@@ 4.2 Solving $\mathbf{Q}_{\text{init}}$ and $\mathbf{K}_{\text{init}}$
There exist numerous approaches to solve Eq. 9 for $\mathbf{Q}_{\text{init}}^{\hphantom{T}}$ and $\mathbf{K}_{\text{init}}^{\hphantom{T}}$ with known $\mathbf{H}_{\text{impulse}}^{\hphantom{T}}$ and $\mathbf{P}$. Here, we apply an SVD-based method. First, we change Eq. 9 to exclude the Softmax function as
@@
With Eq. 11, we do SVD on the right-hand side and get a low-rank approximation of $\mathbf{Q}_{\text{init}}^{\hphantom{T}}$ and $\mathbf{K}_{\text{init}}^{\hphantom{T}}$:
@@
[…]

### v2

@@ 1 Introduction
To bridge this performance gap, several strategies have been proposed. These include self-supervised pretraining on large-scale datasets [dosovitskiy2020vit, touvron2021training], advanced data augmentation techniques [yun2019cutmix, cubuk2020randaugment], and hybrid architectures that incorporate convolutional layers into Vision Transformers [wu2021cvt, liu2021efficient, yuan2021incorporating, li2023uniformer, dai2021coatnet]. More recently, Zhang et al[zhang2022unveiling] explored the use of pretrained weights to initialize ViTs. Building on this idea, subsequent works have shown that carefully designed network initialization strategies can enhance ViT performance on small-scale datasets without modifying the model architecture. In particular, Trockman and Kolter [trockman2023mimetic] introduced mimetic initialization that replicates the weight distribution of pretrained ViTs. Similarly, Xu et al[xu2024initializing] proposed directly sampling weights from large pretrained models.
@@
To overcome the limitations of existing approaches, we propose a novel “structured initialization” strategy for ViTs that imposes convolutional structures on attention maps without requiring any pretrained models. Our approach is grounded in our theoretical insights, aligned with the findings of recent work [cazenavette2025rethinking], that randomly initialized depthwise convolution filters can match the performance of their trained counterparts in models such as ConvMixer [trockman2022patches] and ResNet [he2016deep]. Inspired by this, we develop an initialization strategy for ViTs based on random impulse convolution kernels, which impart locality and structure directly into the attention maps, as shown in Fig.˜1. This structured initialization yields inductive biases characteristic of CNNs, enabling ViTs to generalize effectively on small-scale datasets, while preserving their adaptability for large-scale applications. Unlike prior methods that modify ViT architectures by integrating convolutional components, our method preserves the original transformer structure, making it broadly applicable across a variety of ViT variants.
@@ 2 Related Work
Many efforts have aimed to incorporate a convolutional inductive bias into ViTs through architectural modifications. [dai2021coatnet] proposed to combine convolution and self-attention by mixing the convolutional self-attention layers. [pan2021integration, li2023uniformer]introduced hybrid models wherein the output of each layer is a summation of convolution and self-attention. [wu2021cvt] explored using convolution for token projections within self-attention, while [yuan2021incorporating] showed promising results by inserting a depthwise convolution before the self-attention map. [d2021convit] introduced gated positional self-attention to imply a soft convolution inductive bias. Although these techniques have been proven effective, they aim to introduce the inductive bias of convolution through architectural choices. Our approach, on the other hand, stands out by not requiring any modifications to the architecture, retaining the generalizability to be seamlessly applied to different settings.
@@
The exploration of applying inductive bias through initialization within a transformer is limited to date. [zhang2022unveiling] posited that the benefit of pretrained models in ViTs can be interpreted as a more effective strategy for initialization. [trockman2023mimetic, trockman2022understanding] recently investigated the empirical distributions of self-attention weights, learned from large-scale datasets, and proposed a mimetic initialization strategy. While this approach lies between structured and generative initialization, it relies on the pretraining results of large models. [samragh2023weight, samragh2024scaling, xu2024initializing, li2024surprising] directly sampled weights from pretrained large-scale models as initialization for smaller models. While effective, the sampled weights must follow the distribution of these pretrained weights. A key difference in our approach is that our method does not require offline knowledge of pretrained models. Instead, our initialized structure is derived from a theoretical analysis of convolution layers.
@@
Since their introduction [vaswani2017attention, dosovitskiy2020vit], the relationship between transformers and CNNs has been a topic of immense interest. [andreoli2019convolution] studied the structural similarities between attention and convolution, bridging them into a unified framework. Building on this, [cordonnier2020relationship] demonstrated that self-attention layers can express any convolutional layers through a careful theoretical construction. While these studies highlighted the functional equivalence between self-attention in ViTs and convolutional spatial mixing in CNNs, they did not delve into how the inductive bias of ViTs could be adapted through this theoretical connection. In contrast, our work offers a theoretical insight: a random convolutional impulse filter can be effectively approximated by softmax self-attention.
@@ 3 Why Random Impulse Filters Work?
It is well established that both ConvMixer and ViTs use alternating blocks of spatial and channel mixing, where ViTs replace the spatial convolutions in ConvNets with attention mechanisms. A fundamental difference, except for the receptive field, lies in their parameterization: spatial convolutions with hundreds of channels typically use distinct kernels for each channel, whereas self-attention relies on a shared mechanism with only a limited number ( ${\sim}\,10$) of attention heads. By introducing a key observation that input embeddings are often rank-deficient, we demonstrate that as long as the spatial kernels or attention patterns sufficiently span the kernel space, the uniqueness or repetition of individual kernels becomes less critical. This insight, supported by a recent empirical observation [cazenavette2025rethinking], offers a new perspective on the relationship between ViTs and CNNs, motivating the use of impulse-based structures to embed convolutional inductive biases into the attention map initialization.
In recent work [cazenavette2025rethinking], Cazenavette et al. demonstrated a remarkable performance of randomly initialized convolution filters in ConvMixer and ResNet when solely learning the channel mixing parameters. However, they failed to offer any insights into the underlying reasons. In this section, we provide a theoretical analysis of how solely learning channel mixing can be sufficient for achieving reasonably good performance. Our theoretical findings are significant as they establish a conceptual link between the architecture of ConvMixer and the initialization of ViT, offering a deeper understanding of desired properties for spatial mixing matrices. Without losing generality, we have omitted activations (e.g., GeLU, ReLU, etc.), bias, batch normalization, and skip connections in our equations for clarity.
@@
Let us define the patch embeddings or intermediate layer outputs in ConvMixer as $\mathbf{X}\,{=}\,[\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{D}]$, where $D$ is the number of channels and $N$ is the number of pixels in the vectorized patch $\mathbf{x}\,{\in}\,\mathbb{R}^{N}$. An interesting observation is the rank (stable rank, defined as $\sum\sigma^{2}\,{/}\,\sigma_{max}^{2}$) of $\mathbf{X}$ is consistently much smaller than the minimum dimension $\mbox{min}(N,D)$ of $\mathbf{X}$, indicating a significant amount of redundancy in patch embeddings or intermediate layer outputs in deep networks. This rank deficiency is common in various deep neural networks [feng2022rank], especially in ViTs [noci2022signal, naderi2024mind, ji2025always].
@@
(M–k Spanned Set): Let $\mathcal{V}\,{=}\,\{\mathbf{v}_{1},\mathbf{v}_{2},\ldots,\mathbf{v}_{N}\}\subset\mathbb{R}^{d}$ be a finite set of vectors. We say that $\mathcal{V}$ is MMkk spanned (on $W$) if there exists a partition of $\mathcal{V}$ into at least $k$ non-overlapping subsets:
@@
such that there exist a MM–dimensional subspace $W\subset\mathbb{R}^{d}$ in the span of each subset $\mathcal{V}_{i}$
@@
A ConvMixer block $\mathbf{T}$ consists of a spatial mixing layer $\mathbf{T}_{S}(\>\cdot\>;\mathbf{H})$ with convolution filters $\mathbf{H}$ and a channel mixing layer $\mathbf{T}_{C}$. $\mathbf{T}^{\prime}$ is another ConvMixer block composed of $\mathbf{T}^{\prime}_{S}(\>\cdot\>;\mathbf{H}^{\prime})$ and $\mathbf{T}^{\prime}_{C}$. Let $k$ be the rank of input $\mathbf{X}$. If $\mathbf{H}$ is MMkk spanned on $W$, then for any set of filters $\mathbf{H}^{\prime}\subset W$ and any $\mathbf{T}^{\prime}_{C}$, there always exists a $\mathbf{T}_{C}$ such that $\mathbf{T}(\mathbf{X})\,{=}\,\mathbf{T}^{\prime}(\mathbf{X})$.
For simplicity, we include the full proof in the appendix. Note that since $H$ are convolution matrices, their span lies in $\mathbb{R}^{f{\times}f}$ instead of $\mathbb{R}^{N{\times}N}$, where $f$ is the kernel size. In practice, the fact that $D\,{\geq}\,kf^{2}$ indicates that randomly initialized spatial convolution kernels are $f^{2}$$k$ spanned and satisfy Proposition 1, as illustrated in Fig.˜2. Consequently, any trained results of $\mathbf{T}^{\prime}_{C}$ and $\mathbf{T}^{\prime}_{S}$ can be achieved by solely training the $\mathbf{T}_{C}$, while keeping a fixed spatial mixing layer $\mathbf{T}_{S}$. Hence, the following corollaries can be obtained, and Corollary 1 explains the phenomenon found in [cazenavette2025rethinking], mentioned at the beginning of the section. Corollary 2 inspires our proposed structure initialization for ViTs. The related experimental evidence of these corollaries and our proposition is given in the appendix.
@@
Spatial convolution filters with all ones (referred to as “box” filters) will not perform well since they are 11kk spanned and can produce only averaging values.
@@ 4.1 Expected Initialized Attention Map Structure
ConvMixer and ViT share most of the components in their architectures. The gap in their performance on small-scale datasets stems from their architectural choices regarding the spatial mixing matrix. Although depthwise convolution (ConvMixer) and multi-head self-attention (ViT) may appear distinct at first glance, their underlying goal remains the same: to identify spatial patterns indicated by the spatial mixing matrix. As defined in Sec.˜3, similar to the spatial mixing step in ConvMixer defined in Eq.˜1, the spatial mixing step of multi-head attention can be expressed as
@@
where $d$ represents the feature dimension in each head, typically set to D/hD{/}h, with $h$ being the number of heads, and the matrices $\mathbf{M}_{i}$ for multi-head self-attention can be expressed as follows:
@@
It is worth noting that in Eq.˜1, the spatial matrices $\mathbf{H}$ are in convolutional structure, resulting in a span of $\mathbb{R}^{f{\times}f}$ instead of $\mathbb{R}^{N{\times}N}$, despite each $\mathbf{H}_{i}$ having a size of $N\,{\times}\,N$. This structural constraint ensures that CNNs focus on local features but struggle to capture long-range dependencies. In contrast, the span of spatial matrices $\mathbf{M}$ in Sec.˜4.1 is $\mathbb{R}^{N{\times}N}$, allowing for greater learning capacity without these limitations. However, a randomly initialized $\mathbf{Q}$ and $\mathbf{K}$ contain no structural information, resulting in random matrices as depicted in the bottom left of Fig.˜1.
Leveraging this insight, we propose to initialize the attention map for each head in ViT to a convolutional structure as denoted in the bottom right of Fig.˜1. Our initialization strategy preserves both the advantage of locality and the capacity to learn long-range information. For clarity and brevity, the following discussions will focus only on one head of multi-head self-attention. Therefore, from Sec.˜4.1 and Eq.˜1, our structured initialization strategy can be represented as
$\mathbf{T}^{\text{ViT}}_{S}(\mathbf{X};\mathbf{M})\xleftarrow{\text{init}}\mathbf{T}^{\text{Conv}}_{S}(\mathbf{X};\mathbf{M})\,\Rightarrow\,\mathbf{M}_{\text{init}}^{\hphantom{T}}=\mbox{softmax}(\mathbf{X}\mathbf{Q}_{\text{init}}^{\hphantom{T}}\mathbf{K}_{\text{init}}^{T}\mathbf{X}^{T})\approx\mathbf{H}.$ (8)
@@
Any random convolution filters that satisfy proposition 1 could be a choice of initialization of attention maps to introduce inductive bias. However, random convolution filters $\mathbf{H}$ usually contain both positive and negative values, while the output of the softmax function is always positive, making Eq.˜8 unreachable. One straightforward option is to use random positive convolution filters with a normalized sum of one, following the property of softmax. However, this approach often proves inefficient as the patterns are too complicated for a softmax function to handle with $\mathbf{Q}\mathbf{K}$ being of low rank. In [tarzanagh2023transformers], the authors found that the softmax attention map serves as a feature selection function, which typically tends to select a single related feature. In convolution filters, this selection can be parameterized as impulse filters. According to Corollary 2, random impulse filters are also $f^{2}$$k$ spanned. In conclusion, when initializing a softmax attention map, the most straightforward and suitable choice is random impulse convolution filters.
@@
With the use of impulse filters and the pseudo input, Eq.˜8 becomes
@@ 4.2 Solving $\mathbf{Q}_{\text{init}}$ and $\mathbf{K}_{\text{init}}$
There exist numerous approaches to solve Eq.˜9 for $\mathbf{Q}_{\text{init}}^{\hphantom{T}}$ and $\mathbf{K}_{\text{init}}^{\hphantom{T}}$ with known $\mathbf{H}_{\text{impulse}}^{\hphantom{T}}$ and $\mathbf{P}$. Here, we apply an SVD-based method. First, we change Eq.˜9 to exclude the Softmax function as
@@
With Eq.˜11, we do SVD on the right-hand side and get a low-rank approximation of $\mathbf{Q}_{\text{init}}^{\hphantom{T}}$ and $\mathbf{K}_{\text{init}}^{\hphantom{T}}$:
@@
In ViT models with d= 64d\,{=}\,64, we use a 3× $3\,{\times}\,$3 convolution filters by setting f= 3f\,{=}\,3. We find that the random values in $\mathbf{Z}$ are not decisive for our initialization strategy. To ensure a clear impulse structure, the initialization is primarily governed by selecting an appropriate ratio of $\alpha$ to $\beta$. Empirically, we choose a large ratio such that $\alpha{:}\beta\,{=}\,40{:}1$ with $\gamma\,{=}\,2$. Notably, the exact values of $\alpha$ and $\beta$ are not strictly constrained due to the normalization step, which scales these parameters. The pseudo code for our initialization strategy can be found in the appendix.